I need to solve the following equation via the factoring method, but I don't know where to start. If that information is provided, I will be capable of solving the rest by hand. Thank you.
6x^2+2x-4=0
Once more, thank you one thousand times over.

Answer by addingup(3677)


6x^2 + 2x - 4 = 0
factor the left side:
2(x + 1)(3x - 2) = 0
Divide both sides by 2:
(x + 1)(3x - 2) = 0
x + 1 = 0 or 3x - 2 = 0
I'm sure you can take it from here:
x = -1 or 3x = 2
x = -1 or x = 2/3